OpenStudy (anonymous):

find the surface area of a square pyramid with base edge 80m and slant height 100m

OpenStudy (anonymous):

Sides are triangles of base 80m and height 100m. Area of each is (1/2)BH = (1/2)*80*100=4000m^2. There are four sides, so the total side area is 16000 m^2. The bottom is a square of side 80m, or area 6400 m^2. Therefore the total surface area is 16000+6400 = 22400 m^2.